OpenStudy (anonymous):

how do i find the percent composition for this pb02

OpenStudy (abb0t):

Find the mass of each individual element and divide by the molar mass. then multiply by 100%.

OpenStudy (nincompoop):

I am not sure what pb is and 02 so the answer to your question is unknown. there's not one element in there

thomaster (thomaster):

\(\sf\large PbO_2~=~Lead~dioxide~\ddot\smile\)

thomaster (thomaster):

i think he means that :P

OpenStudy (nincompoop):

that could have been Phosphorus and Boron

thomaster (thomaster):

true :P phosphorus boron oxygen

thomaster (thomaster):

meh we'll see when @dave6heat responds :)

OpenStudy (abb0t):

Don't just assume, @thomaster !! That's so rude!

OpenStudy (anonymous):

um guys it's lead dioxide so whats the percent composition

OpenStudy (abb0t):

Like I said up above: Find the mass of each individual element and divide by the molar mass. then multiply by 100%.

OpenStudy (abb0t):

You may want to take out your periodic table for this.

OpenStudy (anonymous):

yes i have my periodic table thanks

OpenStudy (nincompoop):

then start counting masses, sum them all up and then like what abb0t said compare each element or allotrope against that sum and then multiply it by 100.

OpenStudy (nincompoop):

Pb = 207.2 amu x 1 = 207.2 amu O = 15.999 amu x 2 = 31.998 amu 207.2 + 31.998 = 239.198 \[(207.2 \div 239.198) \times 100 =\] \[(31.998 \div 239.198) \times 100 = \]